OpenStudy (anonymous):

can someone pleaseeee help me factor completely, if polynomial is not factorable write prime A. b^2 -16b +16 B. 10c^2 -17c +3 C. p^2+2p-24 D. c^3 + 8

OpenStudy (anonymous):

B.

OpenStudy (anonymous):

a= prime b=(5c-1)(2c-3) c=(p-4)(p+6) d=(c+2)({c}^{2}-2c+4)